Radovilsky Manufacturing Company, in Hayward, California, makes flashing lights for toys. The company operates its production facility 300 days per year. It has orders for about 12,500 flashing lights per year and has the capability of producing 105 per day. Setting up the light production cast $49. The cost of each light is $0.95. The holding cost is $0.15 per light per year. What is the optimal size of the production run? 
Answer #1

Answer:


Given the data,

D = Annual demand =12500.

O = ordering cost/setup cost =$49,

P= production rate per year = 300

 H = holding cost per year =$0.10

Optimal size of the production run is, Q = Sqrt ((2xOxD)/H))

= Sqrt ((2X49X12500)/0.1) = 3500 units.


answered by: Subrahmanyam golla